OpenStudy (anonymous):

Meredith works at a horse stable. Her job is to exercise the horses. The stable has a circular area that Meredith runs the horses within. The diameter of the fencing is 16 yards. What is the length of the fence? 16π yards 4π yards 16 yards 4 yards

Parth (parthkohli):

Do you know how to find the circumference given the diameter?

Parth (parthkohli):

\(C = \pi \times d\)

OpenStudy (anonymous):

50.24?

Parth (parthkohli):

Don't use a calculator! Solve in terms of pi.

Parth (parthkohli):

What is \(16 \times \pi\)?

OpenStudy (anonymous):

uhm...50.26

Parth (parthkohli):

Don't use a calculator! Just keep it simple!

Parth (parthkohli):

Like \(16 \times x = 16x\), \(16 \times \pi =\) ?

OpenStudy (anonymous):

oh well i get what you mean but i cant type it on here

Parth (parthkohli):

What do you get? just type \(\pi\) as pi.

OpenStudy (anonymous):

16pi

Parth (parthkohli):

That's it!

OpenStudy (anonymous):

so is that the answer?

Parth (parthkohli):

Yuppers

OpenStudy (anonymous):

thank yoou